OpenStudy (anonymous):

Solve for t in lnt - ln (t^2) =5

OpenStudy (anonymous):

how would you use e to cross off ln?

OpenStudy (anonymous):

\[\ln(t) - \ln(t^2) \implies \ln \left( \frac{ t }{ t^2 } \right)\]

OpenStudy (anonymous):

the answer would be t = 1/e^5 correct?

OpenStudy (anonymous):

Yes, that seems good to me

OpenStudy (anonymous):

okay thank you. could you help me with another one?

OpenStudy (anonymous):

solve for x in log2x+log2(x-2)=3

OpenStudy (anonymous):

Pretty much the same process

OpenStudy (anonymous):

x=-4 and x=2?

OpenStudy (anonymous):

\[\log_b(xy) \implies \log_b(x)+\log_b(y)\]

OpenStudy (anonymous):

Idk can you show your work, and I'll see if your steps are right or not?

OpenStudy (anonymous):

yeah one second

OpenStudy (anonymous):

x(x+2)=2^3 x^2+2x=8 x^2+2x-8=0

OpenStudy (anonymous):

It's x-2 i thought

OpenStudy (anonymous):

oh your right

OpenStudy (anonymous):

so would i just switch the negative and positive around

OpenStudy (anonymous):

Yeah, so x = 4 is your solution

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

What about 2^5x/2^x when trying to simplify?

OpenStudy (anonymous):

Check this out, I made this a while back it's a basic tutorial on logarithm stuff, if you get stuck just ask then, give it a read: http://openstudy.com/users/iambatman#/updates/5403fdfde4b0f2ed1e14206a

OpenStudy (anonymous):

would it be 5?

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

\[\frac{ 2^{5x} }{ 2^x } \implies 2^{4x}\]

OpenStudy (anonymous):

which is just 16^x

OpenStudy (anonymous):

oh, is it that you just divide the one 2^x?

OpenStudy (anonymous):

When you divide exponents you subtract, it's like when you're multiplying and you have exponents you add them, right?

OpenStudy (anonymous):

yes